Ln1/e^3=-3

Ln1/e^3=-3
Answer #1

1 3

= In (e-3)

apply log rule loga (r)b loga ()

-3 ln (e

apply log rule \large \log _a\left(a\right)=1

\large =-3(1)

\large {\color{Red} =-3}